# Decentralization of Generative AI via Mixture of Experts for Wireless Networks: A Comprehensive Survey

### v1

@@ I-A Background
[…]
MoE with GenAI [31, 32, 33],
GenAI for wireless Networks [34, 35, 21] , minimum width=7.0cm, minimum width=3.0cm, xshift=-2.0cm, text width=6.0cm, yshift = 0.77cm ] ] ] [Section II. Overview of MoE, minimum width=4.0cm, base=midpoint, yshift = -0.07cm [Definition and Core Components, minimum width=4.0cm, yshift = -0.147cm [Expert network [37, 38, 39, 40], gating mechanism,
and output layer, text width=6.0cm, xshift=-2.0cm, ] ] [Classification of Gating Mechanism, minimum width=4.0cm, [Dense gate [41, 42, 37, 43, 44, 45, 46], sparse gate [47, 48, 49],
hard gate [50, 51], and other gates [52, 53, 54, 55] , text width=6.0cm, xshift=-2.0cm, yshift = 0.14cm] ] [MoE with Generative AI, minimum width=4.0cm,[GANs[56, 57, 58], VAEs [59, 60, 61, 62],
DMs [63, 64, 65], Transformers [66, 67, 68, 69, 70]. , text width=6.0cm, xshift=-2cm, yshift = -0.005cm, base=midpoint, ] ] [MoE with Reinforcement Learning, minimum width=4.0cm, [Value-based RL [71, 72, 73], policy-based RL [74, 75, 76, 77],
multi-agent RL [78, 79, 80], and other RL [81, 82, 83] , text width=6.0cm, xshift=-2.0cm, yshift = 0.14cm] ] ] [Section III.
MoE in Wireless Networks, minimum width=4.0cm, base=midpoint, [Advantages of MoE
in Wireless networks, minimum width=4.0cm, [ Enhance model capacity, computational efficiency,
feature learning, and versatility , text width=6.0cm, xshift=-2cm, yshift = -0.145cm, base=midpoint]] [MoE in Various Wireless
Network Scenarios, minimum width=4.0cm, [Vehicular Networks [84, 85], UAV [86, 87, 88], satellite
[89, 90] and other 6G scenarios [91, 92, 93, 94], text width=6.0cm, xshift=-2cm, yshift = -0.145cm, base=midpoint] ] ] [Section IV. Applications of
MoE in Wireless Networking, minimum width=4.0cm, base=midpoint, [Physical Layer Communication
and Resource Management, minimum width=4.0cm, yshift = 0.142cm, [Channel estimation and prediction [95, 96, 97],
signal processing and communication [98, 99, 100, 61],
radio resource allocation [101, 102, 103] , text width=6.0cm, xshift=-2cm, yshift = 0.00cm, base=midpoint, ]] [Network Optimization
and Security, minimum width=4.0cm, yshift = -0.15cm, [Traffic network management [104, 105, 106],
edge computing and distributed systems [107, 108, 109],
security and anomaly detection [110, 111, 112] , text width=6.0cm, xshift=-2cm, ] ] ] [Section V. Case Study and Dataset, minimum width=4.0cm, [Case Study of MoE with DRL, minimum width=4.0cm, [ MoE is integrated into a diffusion model-based DRL
to demonstrate performance improvements in DRL , text width=6.0cm, xshift=-2cm, base=midpoint, yshift = -0.01cm] ] [Open-Source Datasets, minimum width=4.0cm, [NLP [113, 114, 115], vision [116, 117, 118], multimodal
[119, 120, 121], wireless networks [122, 123, 113, 124] , text width=6.0cm, xshift=-2cm, yshift = 0.14cm ]] ] [Section VI. Future Directions, minimum width=4.0cm, ] [Section VII. Conclusion, minimum width=4.0cm, ] ]
@@ I-B Motivation
The framework of mixture of experts (MoE) has emerged as an effective paradigm to address diverse computational demands and resource constraints for wireless communication systems [27]. The fundamentals of MoE is to decompose a large model into multiple expert modules, each functioning as an independent sub-network specialized for specific inputs [41]. The expert modules are dynamically managed by a gating mechanism that selectively activates only a subset of experts based on the characteristics of the input data, rather than activating all modules simultaneously. Compared with traditional AI, GenAI models require more computational resources for training due to their complex mission objectives and inference procedures. State-of-the-art GenAI models, such as DeepSeek [69], GLaM [48], and Switch Transformer [50], have effectively utilized MoE frameworks to expand model capacity while minimizing computational costs. The dynamic expert selection and sparse activation mechanism of MoE can effectively mitigate computational complexity and improve energy efficiency for GenAI models, enabling resource-constrained wireless infrastructures to enhance scalability and reduce resource consumption in complex wireless task [42].
@@ II-B1 Dense Gating Mechanism
Dense gating has been adopted in multiple early investigations [41, 42, 37], and has more recently been applied to large AI models such as LoRAMoE [43], MoLE [44], and DS-MoE [45]. The dense gating mechanism proves effective in tasks demanding comprehensive expert engagement, especially under high uncertainty or diverse feature distributions [46].
@@ II-B2 Sparse Gating Mechanism
While activating all MoE experts enhances model expressiveness, it also incurs significant computational costs, making it impractical for large-scale applications. Sparse gating mechanism uses the Top-K expert selection approach to limit the number of activated experts per input to $K$, tremendously reducing computational overhead [47]. The gating weight of the $n$-th expert $G(\mathbf{x};\Theta)_{n}$ using the Top-K approach is computed as
$\begin{gathered}G(\mathbf{x};\Theta)_{n}=\operatorname{Softmax}\left(\operatorname{Top-K}\left(G(\mathbf{x};\Theta)\right)\right)_{n},\\ \operatorname{Top-K}\left(G(\mathbf{x};\Theta)\right)_{i}=\begin{cases}G(\mathbf{x};\Theta)_{i},&G(\mathbf{x};\Theta)_{i}\text{ is among the }\\ &\text{top-K values of }G(\mathbf{x};\Theta),{}{}{}\\ -\infty,&\text{otherwise. }\end{cases}\end{gathered}$ (3)
Based on the selective activation of sparse gating mechanism, MoE enables the model to scale efficiently while preserving model capacity, which is particularly important in LLMs where computational efficiency is crucial. Examples of sparse gating implementations include GShard [49] and GLaM [48], which significantly improve scalability by activating only a small subset of experts while maintaining high model performance.
@@ II-B3 Hard Gating Mechanism
$G(\mathbf{x};\Theta)_{n}=\begin{cases}1,&n=\arg\max_{i}G(\mathbf{x};\Theta)_{i},\\ 0,&\text{otherwise. }\end{cases}$ (4)
@@
The hard gating mechanism substantially reduces computational overhead, making it particularly suitable for resource-constrained environments. Despite its simplicity, the hard gating mechanism, as utilized in Switch Transformer [50] and DeepSpeed-MoE [139], has demonstrated competitive performance even when activating only one expert per sample. Nonetheless, the hard gating mechanism may confronted with the challenge of imbalanced expert utilization, where certain experts may be over-utilized while others remain underused. Load-balancing techniques [51] and regularization strategies[140] are essential for ensuring more uniform expert utilization. For example, a router Z-loss is introduced to enable a more uniform output proportion of the gating network, thereby enhancing the stability of model training [141]. Given $B$ input data samples, the Z-loss is formulated as
where $\mathbf{x}_{b}$ is the $b$-th input data and the loss function helps the $G\left(\mathbf{x}_{b};\Theta\right)$ values converge towards a more balanced distribution.
@@ II-B4 Other Variant Gating Mechanisms
Lessons Learned: Gating mechanisms are a critical component of the MoE framework, responsible for regulating expert activation and managing the allocation of computational resources. The choice among dense, sparse, hard, and more advanced adaptive gating strategies involves a fundamental trade-off between expressiveness, efficiency, and specialization. Specifically, the dense gating mechanism ensures full expert participation but incurs substantial computational overhead. In contrast, sparse and hard gating mechanims improve efficiency by limiting expert activation, while potentially posing challenges related to stability and capacity utilization. Emerging designs such as multi-gate and hierarchical gating mechanism provide enhanced flexibility by enabling task-aware or multi-tier expert selection, thereby supporting fine-grained specialization. The effectiveness of a gating mechanism depends on balancing computational cost and model capacity, as well as adapting to task complexity and data heterogeneity. As such, the design and deployment of gating mechanisms are pivotal for scaling MoE frameworks to meet the diverse demands of machine learning applications.
@@ II-C MoE with Generative AI
  Diffusion   Models [63] Text-to-image Generation Train specialized denoisers for different denoising intervals to align text and image fidelity Fixed parameters across denoising steps hinder adaptability Employ specialized experts for different denoising stages, adapting to distinct noise characteristics at each time step Sparse experts reduce inference cost while maintaining generation quality
@@
Transformers [68] Language Generation Construct attention and MLP-based MoE for instructed LLMs with post-training Substantial computational requirements and limited adaptability to diverse modalitie Replace FFNs with MoE layers to reduce computation and maintain model capacity Enable experts to handle inputs across language, vision, and multimodal domain
@@
The employment of MoE framework in GenAI models has garnered increasing attention, leveraging MoE experts with generative capabilities to enhance data synthesis, representation learning, and adaptive processing. Recent advancements have demonstrated the effectiveness of MoE in improving the performance of GenAI models, including generative adversarial networks (GANs), variational autoencoders (VAEs), diffusion models (DMs), and Transformer-based architectures [79]. This section explores the integration of MoE and GenAI, highlighting key advancements in optimizing generative performance across diverse applications. Figure 3(c) illustrates an MoE-enhanced GenAI framework based on VAE, and a summary of related approaches is provided in Table II-C.
II-C1 Generative Adversarial Networks
II-C2 Variational Autoencoders
II-C3 Diffusion Models
II-C4 Transformers

### v2

# Mixture of Experts for Decentralized Generative AI and Reinforcement Learning in Wireless Networks: A Comprehensive Survey

@@ I-B Motivation
The framework of mixture of experts (MoE) has emerged as an effective paradigm to address diverse computational demands and resource constraints for wireless communication systems [27]. The fundamentals of MoE are to decompose a large model into multiple expert modules, each functioning as an independent sub-network specialized for specific inputs [41]. The expert modules are dynamically managed by a gating mechanism that selectively activates only a subset of experts based on the characteristics of the input data, rather than activating all modules simultaneously. Compared with traditional AI, GenAI models require more computational resources for training due to their complex mission objectives and inference procedures. State-of-the-art GenAI models, such as DeepSeek [69], GLaM [48], and Switch Transformer [50], have effectively utilized MoE frameworks to expand model capacity while minimizing computational costs. The dynamic expert selection and sparse activation mechanism of MoE can effectively mitigate computational complexity and improve energy efficiency for GenAI models, enabling resource-constrained wireless infrastructures to enhance scalability and reduce resource consumption in complex wireless task [42].
@@ II-B1 Dense Gating Mechanism
Dense gating has been adopted in multiple early investigations [41, 42, 37], and has more recently been applied to large AI models such as LoRAMoE [43], MoLE [44], and DS-MoE [45]. Since the dense gating mechanism activates all experts for each input instance, they have proved effective in tasks demanding comprehensive expert engagement, especially under high uncertainty or diverse feature distributions [46]. However, every expert receives gradients during back-propagation, causing the total computational cost and memory usage to scale linearly with the number of experts $N$. While this scheme simplifies expert selection and enhances model expressiveness, it usually incurs substantial inter-expert communication overhead and computational costs, making it impractical for large-scale applications.[139].
@@ II-B2 Sparse Gating Mechanism
Sparse gating mechanism uses the Top-K expert selection approach to limit the number of activated experts per input to $K\ll N$, which tremendously reduces the computational overhead [47]. The gating weight of the $n$-th expert $G(\mathbf{x};\Theta)_{n}$ using the Top-K approach is computed as
[…]
@@ II-B3 Hard Gating Mechanism
[…]
@@
The hard gating mechanism substantially reduces computational overhead, making it particularly suitable for resource-constrained environments. Despite its simplicity, the hard gating mechanism, as utilized in Switch Transformer [50] and DeepSpeed-MoE [140], has demonstrated competitive performance even when activating only one expert per sample. Nonetheless, the hard gating mechanism may confronted with the challenge of imbalanced expert utilization, where certain experts may be over-utilized while others remain underused, potentially limiting the model’s capacity utilization and expressive generalization.
@@ II-B4 Other Variant Gating Mechanisms
Mitigation of Expert Collapse and Overfitting: Although gating mechanisms effectively regulate expert activation and manage computational resource allocation, issues such as expert collapse and model overfitting persist and require further mitigation, particularly under limited training data. Expert collapse occurs when a small number of experts dominate training, leading to under-utilization of other experts, thereby diminishing model diversity and capacity. Several strategies have been proposed to address the challenge of expert collapse. Load-balancing losses, such as auxiliary entropy-based losses, encourage a balanced distribution of expert activations [51]. For example, a router Z-loss is introduced to enable a more uniform output proportion of the gating network [141]. Given $B$ input data samples, the Z-loss is formulated as
where $\mathbf{x}_{b}$ is the $b$-th input data and the loss function helps the $G\left(\mathbf{x}_{b};\Theta\right)$ values converge towards a more balanced distribution. The load-balancing strategies effectively mitigating expert collapse by penalizing excessive reliance on individual experts. thereby enhancing the stability of model training. Additionally, overfitting problem arises when experts overly adapt to limited training data, reducing their ability to generalize effectively. Regularization methods such as dropout, early stopping, and weight decay are commonly integrated into MoE frameworks to prevent overfitting [142]. More recently, knowledge distillation [143] and expert pruning techniques [144] have emerged as effective approaches to maintain expert diversity, especially under data-scarce conditions. These mechanisms collectively enhance the robustness of MoE, ensuring better generalization across diverse tasks and environments.
Lessons Learned: Gating mechanisms serve as a critical component of the MoE framework, responsible for regulating expert activation and managing the allocation of computational resources. These mechanisms differ significantly in their trade-offs among computational efficiency, representational capacity, and system complexity. Specifically, dense gating activates all experts for every input, ensuring full expert participation and rich expressiveness, but at the cost of substantial computational and memory overhead. Sparse gating, such as Top-k selection, reduces training FLOPs and memory consumption by activating only a subset of experts per input, which is widely adopted in large-scale models such as GShard and GLaM. However, it usually incurs a slight reduction in model capacity and requires effective load-balancing strategies to mitigate expert underutilization. Hard gating deterministically selects a single expert for each input, achieving maximal computational efficiency and making it suitable for resource-constrained scenarios. Nonetheless, hard gating introduces high gradient variance and potential expert imbalance that impairs the stability of the training process. Advanced gating strategies, such as multi-gate and hierarchical gating, support task-specific or multi-level expert selection, enabling finer specialization and better adaptation to heterogeneous inputs, but also increase architectural and training complexity. Overall, each gating strategy offers distinct advantages depending on the target deployment scenarios and resource constraints. As such, the design and deployment of gating mechanisms are pivotal for scaling MoE frameworks to meet the diverse requirements of modern machine learning applications, particularly in dynamic and resource-sensitive environments such as wireless networks.
@@ II-C MoE with Generative AI
The employment of MoE framework in GenAI models has garnered increasing attention, leveraging MoE experts with generative capabilities to enhance data synthesis, representation learning, and adaptive processing. Recent advancements have demonstrated the effectiveness of MoE in improving the performance of GenAI models, including generative adversarial networks (GANs), variational autoencoders (VAEs), diffusion models (DMs), and Transformer-based architectures [79]. This section explores the integration of MoE and GenAI, highlighting key advancements in optimizing generative performance across diverse applications. Figure 3(c) illustrates an MoE-enhanced GenAI framework based on VAE, and a summary of related approaches is provided in Table II.
II-C1 Generative Adversarial Networks
II-C2 Variational Autoencoders
II-C3 Diffusion Models
II-C4 Transformers
@@
  Diffusion   Models [63] Text-to-image Generation Train specialized denoisers for different denoising intervals to align text and image fidelity
Fixed parameters across denoising steps hinder adaptability
  • Employ specialized experts for different denoising stages, adapting to distinct noise characteristics at each time step
  • Sparse experts reduce inference cost while maintaining generation quality
@@
    Transformers [68] Language Generation Construct attention and MLP-based MoE for instructed LLMs with post-training
    Substantial computational requirements and limited adaptability to diverse modalitie
  • Replace FFNs with MoE layers to reduce computation and maintain model capacity
  • Enable experts to handle inputs across language, vision, and multimodal domain